OpenStudy (anonymous):

sqrt((25 tan x )^2 +25). should the answer equate to sec^2 x? or 5 sec x ?

OpenStudy (jamesj):

Well, I assume you mean this \[ \sqrt{25\tan^2 x + 25} \] which is equal to \[5 \sqrt{\tan^2 x + 1} \] because \( 5 = \sqrt{25} \). Can you finish the problem now yourself?

OpenStudy (phi):

I suppose it is worth memorizing the fact that \[ tan^2 x+ 1 = sec^2 x \] though it is definitely worth knowing that \[ sin^2 x + cos^2 x =1 \] we can use this to show the first identity. Of course, you know \[ tan (x) = \frac{sin (x)}{cos( x)} \] if we add 1 to tan^2(x): \[ \frac{sin^2 (x)}{cos^2( x)} + \frac{cos^2(x)}{cos^2(x)} \] which is the same as \[ \frac{sin^2 (x)+cos^2( x)}{cos^2( x)} \] now use the very important identity to make the top 1. We get \[ \frac{1}{cos^2( x)} = sec^2(x) \] so we now know \[ tan^2 x+ 1 = sec^2 x \]

OpenStudy (anonymous):

oh okay got it now thank you